4) You are deciding whether to buy or lease a $30,000 automobile. A three year lease will cost $6,000 per year (paid at the eHi the question is should you buy or lease the automobile? Please ignore the subsequesnt decription. Thank you!
Answer #1

The decision of whether to lease or purchase outright can be made by analysing the present value of the cash flows.

Present value of Cost of leasing:

[{-6000/(1.1)^1} + {-6000/(1.1)^2} + {-6000/(1.1)^3}] = [(-5454.5454) + (-4958.6777) + (-4507.8888)] = $ -14921.1119 (negative is because it is a cash outflow)

Present value of Cost of buying:

[{-Initial cost of purchase} + discounted value at end of 3 years] = [-30000 + {20000/(1.1)^3}] = -30000 + 15026.2960 = $ -14973.7040(negative because it is net cash outflow)

Since the cost of buying is more than the cost of leasing, it can be decided to lease the automobile.